OpenStudy (anonymous):

3x + 2(4 + 6x) = 113

OpenStudy (anonymous):

3x+2(4+6x)=113 3x+8+12x=113 15x=113-8 15x=105 x=105/15 x=7

OpenStudy (anonymous):

3x+8+12x=113 3x+12x+8=113 15x=113-8 15x=105 x=105/15 x=7

OpenStudy (anonymous):

thanks you :)

OpenStudy (anonymous):

Your welcome :)